Simco Corporation bond has a current price of $62,321.30 and will pay the holder of the instrument $100,000 ten years in the future. What is its yield?
Answer #1

Ans 4.84%

FV = Future Value
PV = Present Value
r = rate of interest
n= no of period
FV/ PV = (1 + r )^n
100000 / 62321.30 = (1 + r%)^10
1.604587837545 = (1 + r%)^10
r = 4.84%